2. A companys flesxible budget for 60.000 units of production showed sales of $96,000, variable costs of S36,000, and fixed costs of $26,000. What operating income would be expected if the company produces and sells 70,000 units? Use a contribution margin format. You must show how you calculated each number for credit. Use the template below for all of the remaining problems. Check: Operating income should be greater than $43,000. (3 points) Sales Variable costs Contribution margin Fixed costs Operating income
Answer #1

Sales price per unit=96000/60000=$1.6 per unit

Variable cost per unit=36000/60000=$0.6 per unit

Sales(1.6*70000) 112000
Variable costs(0.6*70000) 42000
Contribution margin 70000
Fixed costs $26000
Operating income $44000.

NOTE:Total fixed costs and variable costs/unit do not change with change in units.